OpenStudy (anonymous):

evaluate the following integral: 2sinh(tan2(theta))sec^2(2(theta)) d(theta) with limits of integration: (lower) -pi/8 , and (upper) pi/8 ..iv partially solved it with a u substitution but I dont know how to convert the limits of integration.. :/

OpenStudy (anonymous):

so far i have 2cosh(u) .. with u = tan2(theta) ... soo it would be: 2cosh(tan(2(theta))) .. but idk what to do with the old limits of integration.. of (-pi/8 and pi/8) ???

OpenStudy (anonymous):

Well you don't want to use tan^2(theta) for u.

OpenStudy (anonymous):

its tan(2theta) not... tan^2(theta) ...

OpenStudy (anonymous):

And is your integral: \[2 \int\limits \sinh(\tan(2\theta))\sec^2(2 \theta)dtheta\]?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

yes... I got to the part : 2cosh(tan(2(pi/8))) - (2cosh(tan(2(-pi/8))))

OpenStudy (anonymous):

Well you need to evaluate: \[\tan(\pm 2 * \frac{\pi}{8})\] For the positive case you just have: \[\tan(\frac{\pi}{4}) = \frac{\sin\left(\frac{\pi}{4} \right)}{\cos \left( \frac{\pi}{4} \right)} = \frac{\frac{1}{\sqrt2}}{\frac{1}{\sqrt2}}=1\] And since tangent is an odd function this means: \[\tan(-x) = -\tan(x); \implies \tan(- \frac{\pi}{4}) = -\tan(\frac{\pi}{4}) = -1\]

OpenStudy (anonymous):

i cant understand your syntax... :/

OpenStudy (anonymous):

Then the integration is easy. And if you want to simplify it even more replace cosh(u) (once you integrate) with: \[\cosh(x) = \frac{e^x + e^{-x}}{2}\]

OpenStudy (anonymous):

Your browser must not be updated if you can't see the latex properly.

OpenStudy (anonymous):

im using internet explorer... :/

OpenStudy (anonymous):

eff.. the answer is zero.. nvm.. thanks though

OpenStudy (anonymous):

Ah, this makes sense. Sinh(x) is an odd function which means that sinh(-x) = -sinh(x) so when you evaluate it over symmetric bounds it is zero.

OpenStudy (anonymous):

thank you! :) knowledge!

OpenStudy (anonymous):

while you're still here... what is the integral of cosh^2?